# Security Analysis of Web Applications Based on Gruyere

### v1

@@ 3.1.4. Application-Level Expression Injection
Expression Language (EL) allows dynamic evaluation of variables, function calls, and logical operations within templates or code. EL injection occurs when user input is directly incorporated into EL expressions without proper validation, enabling attackers to execute arbitrary EL code. This vulnerability typically arises when the server supports EL evaluation but fails to properly handle user input. EL injection can manifest in multiple contexts. In JSP and Java-based environments such as Servlet and Spring MVC, attackers can manipulate EL expressions through unvalidated input to perform unauthorized operations or access sensitive information. In Spring-based applications, the Spring Expression Language (SpEL) supports reflection and Bean manipulation, and improperly handled input may allow attackers to read files or execute system commands. In Apache Struts2, Object-Graph Navigation Language (OGNL) expressions can be exploited to execute malicious Java code. Similarly, in Freemarker templates, unfiltered user input may be leveraged to execute system commands or other unauthorized operations(paulson2005building, ). Effective mitigation strategies include disabling EL evaluation when not required, using secure APIs to handle user input and avoid direct expression parsing, applying parameterized input and whitelisting to prevent direct concatenation of user data into EL expressions, and restricting Java reflection and class access to prevent EL from executing system commands through reflection.
Object-Relational Mapping allows developers to interact with databases using an object-oriented programming paradigm without directly writing SQL statements. However, when user input is passed to ORM queries without proper validation, attackers can craft malicious input to bypass authentication, access sensitive data, or even execute arbitrary SQL commands. This type of vulnerability is referred to as ORM injection. The underlying principles of ORM injection are similar to those of other injection attacks. Effective mitigation primarily involves the use of parameterized queries and rigorous validation of all user inputs to prevent unauthorized database operations(gupta2016enhanced, ).
@@ 3.2. Cross-Site Scripting
Cross-Site Scripting is a type of client-side code injection attack that arises when web applications fail to properly validate or encode user input. In an XSS attack, malicious scripts are injected into legitimate web pages and executed by the end-user’s browser, potentially compromising sensitive information. Commonly affected components include search forms, login interfaces, and other input-handling modules, with the primary target being client-side users. Depending on the browser type and version, the consequences of XSS attacks may vary, including credential theft, cookie capture, session hijacking, identity spoofing, generation of fake traffic, pop-up advertisements, and malware propagation(parameshwaran2015dexterjs, ). Cross-Site Scripting attacks are typically divided into three categories.
The first type is Reflected XSS. In this case, the attacker injects a malicious script into a URL. When the victim clicks the malicious link, the server reflects the injected code in the response, which is then executed in the browser. As shown in Listing 1, the PHP program does not filter user input, which introduces a potential vulnerability.An attacker may construct a malicious URL such as http://example.com/?name=<script>alert(’XSS’)</script> which, when visited, triggers the execution of the injected script in the browser. When a victim visits this link, the script is executed. In a more severe case, an attacker may steal user credentials, as illustrated in Listing 2. The second type is Stored XSS. In this case, the malicious script is permanently stored in a database or file system and is executed whenever a user loads the affected page(zhang2025risk, ). As shown in Listing 3, the program directly stores user input into the database and echoes it without sanitization, leading to a potential security risk. If an attacker submits the payload shown in Listing 4, every user visiting the page will trigger script execution, potentially leading to cookie theft or account compromise. The third type is DOM-Based XSS. This attack occurs entirely on the client side. The attacker manipulates the Document Object Model (DOM) so that malicious scripts are executed in the browser. As shown in Listing 5, the program directly inserts input into the DOM without sanitization, which allows arbitrary script execution. If the victim visits a URL such as http://example.com/?q=<script>alert(’XSS’)</script> the malicious code will be executed in the browser(shen2025blockchain, ).
@@ 3.3. Sensitive Data Exposure
Sensitive data exposure arises primarily from inadequacies in application security, which prevent effective protection of confidential information and provide opportunities for attackers to gain unauthorized access to critical data. Such sensitive information typically includes user identity data (e.g., name, address, national ID, passport number), account credentials (e.g., username, password, tokens), financial information (e.g., credit card numbers, bank accounts), medical records (e.g., medical history, diagnoses), and business data (e.g., API keys, encryption keys, database backups). Common scenarios that may result in sensitive data exposure can be summarized as follows(wang2025ai, ). One situation is the storage of passwords in plaintext. If a database is breached, attackers can directly obtain all user passwords. A secure approach is to store passwords with strong hashing algorithms. Another situation arises when sensitive data is transmitted without encryption. Attackers may intercept the communication and capture account information or passwords. This risk can be reduced by enforcing HTTPS or adopting other secure transmission protocols. Sensitive information can also be leaked through API responses. For instance, if an API returns user passwords or credit card details, attackers may exploit this weakness. A proper defense includes removing sensitive fields from responses, applying strict access control, and minimizing the data revealed through APIs. Data exposure may further occur through code or logs that disclose database usernames, server details, or similar information, which can then be leveraged for brute-force or targeted attacks. Sensitive details should not appear in logs or source code. Finally, misconfigured web servers may allow directory or file exposure, giving attackers access to files such as .git/, backup.sql, config.php, or admin/config.json. Adequate access control and server configuration are necessary to prevent such problems(li2021hybrid, ).
@@ 3.4. Authentication and Access Control Vulnerabilities
Broken authentication refers to security vulnerabilities within the authentication process that may allow attackers to bypass verification steps, impersonate legitimate users, or illegally gain account access. Common risks include weak passwords, credential leaks, improper session management, lack of multi-factor authentication, and insufficient protection against brute-force attacks. Users often choose simple passwords like ”123456”, ”password”, or ”admin”, which are vulnerable to dictionary attacks. Once database credentials are compromised, attackers may leverage this information for credential stuffing attacks across other platforms. Key defenses against such threats include enforcing strict password policies, limiting consecutive failed login attempts, and enabling multi-factor authentication. Session fixation attacks involve attackers controlling session IDs and forcing victims to use them, thereby hijacking identities. Session hijacking attacks rely on XSS or traffic interception to steal session IDs. Mitigating these risks requires combining multiple security strategies, such as using HttpOnly cookies, enabling CSRF protection, and strengthening session management to safeguard user accounts(peng2025mining, ).
Access control is a fundamental security mechanism designed to ensure that users can only access data or perform operations within their authorized scope. When access control is inadequately designed or implemented, attackers may bypass permission restrictions to access, modify, or delete resources without authorization(xiang2025security, ). Common access control vulnerabilities include horizontal privilege escalation, where ordinary users access other users’ data, vertical privilege escalation, where ordinary users gain access to administrative functions, forced browsing by modifying request parameters to reach restricted resources, improper API access control where endpoints lack adequate authentication or authorization, and misconfigured cross-origin resource sharing that may expose sensitive information. Defensive measures involve adhering to the principle of least privilege to grant users only the access necessary to perform their tasks, enforcing strict permission checks on the server side instead of relying on frontend validation, establishing a robust role-based access control system, ensuring that APIs require authentication and authorization for all endpoints, and employing unpredictable resource identifiers while validating access rights on the server side to prevent unauthorized access(li2025interaction, ).
@@ 3.5. Vulnerabilities in Data Parsing and Deserialization
XML External Entity (XXE) injection is caused by flaws in the XML parsing process. Attackers may exploit the parser’s handling of external entities, which can lead to sensitive data disclosure, local file access, server-side request forgery, or denial-of-service (DoS) attacks. XXE attacks are generally divided into four categories. The first type is Local File Disclosure. Attackers may design a malicious XML to retrieve sensitive files from the server. As shown in Listing 6, the entity definition references the system file /etc/passwd, and the parser replaces the entity with the file contents. The second type is Server-Side Request Forgery (SSRF)(niu2025natlm, ). When external URLs are allowed, attackers may induce the server to send unauthorized requests. Listing 7 shows an example where the parser attempts to access an internal administrative resource. The third type is Remote File Inclusion. If the parser accepts remote resources, attackers can supply malicious files. As shown in Listing 8, the entity loads a crafted XML from an attacker-controlled domain, which may lead to execution of malicious code. The fourth type is Blind XXE. Even without direct output, attackers can exfiltrate data through external channels. Listing 9 illustrates how a crafted request leaks file content via DNS or HTTP queries(zhou2025blockchain, ).
@@
XXE is a widespread vulnerability in web applications and must be mitigated by disabling external entity resolution, using secure parsers, and enforcing strict input validation. Insecure deserialization occurs when applications process user-supplied serialized objects without validation. This vulnerability generally falls into three categories. The first type is Arbitrary Code Execution. Attackers craft malicious objects that trigger code execution during deserialization. As shown in Listing 10, Java’s ObjectInputStream.readObject() can be abused to execute attacker-supplied code. The second type is Logical Attacks. Attackers tamper with serialized data to bypass authentication or access control. Listing 11 shows an example where PHP’s unserialize() may be used to escalate privileges. The third type is Denial of Service. Attackers send large or recursive malicious serialized data to exhaust system resources. As shown in Listing 12, Python’s pickle.loads() may trigger resource consumption. Insecure deserialization is a critical vulnerability that can result in remote code execution, privilege escalation, or denial of service. Mitigation requires disabling deserialization of untrusted data, validating and sanitizing all inputs, applying integrity checks or digital signatures, and deploying web Application Firewalls(jin2025blockchain, ).

### v2

@@ 3.1.4. Application-Level Expression Injection
Expression Language (EL) allows dynamic evaluation of variables, function calls, and logical operations within templates or code. EL injection occurs when user input is directly incorporated into EL expressions without proper validation, enabling attackers to execute arbitrary EL code. This vulnerability typically arises when the server supports EL evaluation but fails to properly handle user input. EL injection can manifest in multiple contexts. In JSP and Java-based environments such as Servlet and Spring MVC, attackers can manipulate EL expressions through unvalidated input to perform unauthorized operations or access sensitive information. In Spring-based applications, the Spring Expression Language (SpEL) supports reflection and Bean manipulation, and improperly handled input may allow attackers to read files or execute system commands. In Apache Struts2, Object-Graph Navigation Language (OGNL) expressions can be exploited to execute malicious Java code. Similarly, in Freemarker templates, unfiltered user input may be leveraged to execute system commands or other unauthorized operations(Paulson, 2005). Effective mitigation strategies include disabling EL evaluation when not required, using secure APIs to handle user input and avoid direct expression parsing, applying parameterized input and whitelisting to prevent direct concatenation of user data into EL expressions, and restricting Java reflection and class access to prevent EL from executing system commands through reflection.
Object-Relational Mapping allows developers to interact with databases using an object-oriented programming paradigm without directly writing SQL statements. However, when user input is passed to ORM queries without proper validation, attackers can craft malicious input to bypass authentication, access sensitive data, or even execute arbitrary SQL commands. This type of vulnerability is referred to as ORM injection. The underlying principles of ORM injection are similar to those of other injection attacks. Effective mitigation primarily involves the use of parameterized queries and rigorous validation of all user inputs to prevent unauthorized database operations(Gupta and Gupta, 2016).
@@ 3.2. Cross-Site Scripting
Cross-Site Scripting is a type of client-side code injection attack that arises when web applications fail to properly validate or encode user input. In an XSS attack, malicious scripts are injected into legitimate web pages and executed by the end-user’s browser, potentially compromising sensitive information. Commonly affected components include search forms, login interfaces, and other input-handling modules, with the primary target being client-side users. Depending on the browser type and version, the consequences of XSS attacks may vary, including credential theft, cookie capture, session hijacking, identity spoofing, generation of fake traffic, pop-up advertisements, and malware propagation(Parameshwaran et al., 2015). Cross-Site Scripting attacks are typically divided into three categories.
The first type is Reflected XSS. In this case, the attacker injects a malicious script into a URL. When the victim clicks the malicious link, the server reflects the injected code in the response, which is then executed in the browser. As shown in Listing 1, the PHP program does not filter user input, which introduces a potential vulnerability.An attacker may construct a malicious URL such as http://example.com/?name=<script>alert(’XSS’)</script> which, when visited, triggers the execution of the injected script in the browser. When a victim visits this link, the script is executed. In a more severe case, an attacker may steal user credentials, as illustrated in Listing 2. The second type is Stored XSS. In this case, the malicious script is permanently stored in a database or file system and is executed whenever a user loads the affected page(Zhang et al., 2025b). As shown in Listing 3, the program directly stores user input into the database and echoes it without sanitization, leading to a potential security risk. If an attacker submits the payload shown in Listing 4, every user visiting the page will trigger script execution, potentially leading to cookie theft or account compromise. The third type is DOM-Based XSS. This attack occurs entirely on the client side. The attacker manipulates the Document Object Model (DOM) so that malicious scripts are executed in the browser. As shown in Listing 5, the program directly inserts input into the DOM without sanitization, which allows arbitrary script execution. If the victim visits a URL such as http://example.com/?q=<script>alert(’XSS’)</script> the malicious code will be executed in the browser(Shen et al., 2025).
@@ 3.3. Sensitive Data Exposure
Sensitive data exposure arises primarily from inadequacies in application security, which prevent effective protection of confidential information and provide opportunities for attackers to gain unauthorized access to critical data. Such sensitive information typically includes user identity data (e.g., name, address, national ID, passport number), account credentials (e.g., username, password, tokens), financial information (e.g., credit card numbers, bank accounts), medical records (e.g., medical history, diagnoses), and business data (e.g., API keys, encryption keys, database backups). Common scenarios that may result in sensitive data exposure can be summarized as follows(Wang and Li, 2025). One situation is the storage of passwords in plaintext. If a database is breached, attackers can directly obtain all user passwords. A secure approach is to store passwords with strong hashing algorithms. Another situation arises when sensitive data is transmitted without encryption. Attackers may intercept the communication and capture account information or passwords. This risk can be reduced by enforcing HTTPS or adopting other secure transmission protocols. Sensitive information can also be leaked through API responses. For instance, if an API returns user passwords or credit card details, attackers may exploit this weakness. A proper defense includes removing sensitive fields from responses, applying strict access control, and minimizing the data revealed through APIs. Data exposure may further occur through code or logs that disclose database usernames, server details, or similar information, which can then be leveraged for brute-force or targeted attacks. Sensitive details should not appear in logs or source code. Finally, misconfigured web servers may allow directory or file exposure, giving attackers access to files such as .git/, backup.sql, config.php, or admin/config.json. Adequate access control and server configuration are necessary to prevent such problems(Li and others, 2021).
@@ 3.4. Authentication and Access Control Vulnerabilities
Broken authentication refers to security vulnerabilities within the authentication process that may allow attackers to bypass verification steps, impersonate legitimate users, or illegally gain account access. Common risks include weak passwords, credential leaks, improper session management, lack of multi-factor authentication, and insufficient protection against brute-force attacks. Users often choose simple passwords like ”123456”, ”password”, or ”admin”, which are vulnerable to dictionary attacks. Once database credentials are compromised, attackers may leverage this information for credential stuffing attacks across other platforms. Key defenses against such threats include enforcing strict password policies, limiting consecutive failed login attempts, and enabling multi-factor authentication. Session fixation attacks involve attackers controlling session IDs and forcing victims to use them, thereby hijacking identities. Session hijacking attacks rely on XSS or traffic interception to steal session IDs. Mitigating these risks requires combining multiple security strategies, such as using HttpOnly cookies, enabling CSRF protection, and strengthening session management to safeguard user accounts(Peng et al., 2025a).
Access control is a fundamental security mechanism designed to ensure that users can only access data or perform operations within their authorized scope. When access control is inadequately designed or implemented, attackers may bypass permission restrictions to access, modify, or delete resources without authorization(Xiang et al., 2025). Common access control vulnerabilities include horizontal privilege escalation, where ordinary users access other users’ data, vertical privilege escalation, where ordinary users gain access to administrative functions, forced browsing by modifying request parameters to reach restricted resources, improper API access control where endpoints lack adequate authentication or authorization, and misconfigured cross-origin resource sharing that may expose sensitive information. Defensive measures involve adhering to the principle of least privilege to grant users only the access necessary to perform their tasks, enforcing strict permission checks on the server side instead of relying on frontend validation, establishing a robust role-based access control system, ensuring that APIs require authentication and authorization for all endpoints, and employing unpredictable resource identifiers while validating access rights on the server side to prevent unauthorized access(Li et al., 2025).
@@ 3.5. Vulnerabilities in Data Parsing and Deserialization
XML External Entity (XXE) injection is caused by flaws in the XML parsing process. Attackers may exploit the parser’s handling of external entities, which can lead to sensitive data disclosure, local file access, server-side request forgery, or denial-of-service (DoS) attacks. XXE attacks are generally divided into four categories. The first type is Local File Disclosure. Attackers may design a malicious XML to retrieve sensitive files from the server. As shown in Listing 6, the entity definition references the system file /etc/passwd, and the parser replaces the entity with the file contents. The second type is Server-Side Request Forgery (SSRF)(Niu et al., 2025). When external URLs are allowed, attackers may induce the server to send unauthorized requests. Listing 7 shows an example where the parser attempts to access an internal administrative resource. The third type is Remote File Inclusion. If the parser accepts remote resources, attackers can supply malicious files. As shown in Listing 8, the entity loads a crafted XML from an attacker-controlled domain, which may lead to execution of malicious code. The fourth type is Blind XXE. Even without direct output, attackers can exfiltrate data through external channels. Listing 9 illustrates how a crafted request leaks file content via DNS or HTTP queries(Zhou et al., 2025).
@@
XXE is a widespread vulnerability in web applications and must be mitigated by disabling external entity resolution, using secure parsers, and enforcing strict input validation. Insecure deserialization occurs when applications process user-supplied serialized objects without validation. This vulnerability generally falls into three categories. The first type is Arbitrary Code Execution. Attackers craft malicious objects that trigger code execution during deserialization. As shown in Listing 10, Java’s ObjectInputStream.readObject() can be abused to execute attacker-supplied code. The second type is Logical Attacks. Attackers tamper with serialized data to bypass authentication or access control. Listing 11 shows an example where PHP’s unserialize() may be used to escalate privileges. The third type is Denial of Service. Attackers send large or recursive malicious serialized data to exhaust system resources. As shown in Listing 12, Python’s pickle.loads() may trigger resource consumption. Insecure deserialization is a critical vulnerability that can result in remote code execution, privilege escalation, or denial of service. Mitigation requires disabling deserialization of untrusted data, validating and sanitizing all inputs, applying integrity checks or digital signatures, and deploying web Application Firewalls(Jin et al., 2025).